# Depends-Kotlin: A Cross-Language Kotlin Dependency Extractor

### v1

@@ 3.2. Performance Evaluation
In order to evaluate Depends-Kotlin’s performance, we selected three Kotlin projects of different sizes (RubberPicker [^2] https://github.com/Chrisvin/RubberPicker (Kotlin code only), sqlex [^3] https://github.com/sqlex/sqlex (Kotlin and Java code), and ktor [^4] https://github.com/ktorio/ktor (Kotlin code only)) for analysis. We ran Depends-Kotlin on these three projects with a computer (AMD Ryzen 7 4800H@2.9GHz,8GB RAM), and the results of the analysis by Depends-Kotlin are shown in Table  3.
Table 3. Entity and relations extracted in 3 projects.
Entity RubberPicker sqlex ktor
Package 2 45 255
File 11 231 1922
Class, Interface 16 281 3555
Method Member 139 897 13740
Java Field - 162 -
Kotlin Property 71 133 3040
Dependency Relation
Import 7 300 60
Extend 0 45 476
Call 24 711 2786
Use 74 698 4237
Parameter 8 162 1103
Return 2 131 464
Extension 0 45 390
Delegate 0 0 18
JDependsOnK - 1 -
KDependsOnJ - 343 -
#Total_Relation 115 2436 9534
Consumed Time 8s 39s 581s
From this table, we can observe that in a small Kotlin project like RubberPicker, there are no delegate or extension dependency relations. In contrast, delegate or extension relationships can be identified in a larger Kotlin project such as ktor. In the sqlex project, which includes both Java and Kotlin code, we can find the KDependsOnJ and JDependsOnK dependency relations, indicating Kotlin and Java’s interactions.
Table 3 also shows that, for the RubberPicker project with 16 classes, 139 methods and 115 dependency relations, Depends-Kotlin can resolve it in 8s. For the ktor project with 3,555 classes and 9,534 dependency relations, Depends-Kotlin requires 581 seconds for resolution. For a medium project sqlex with 281 classes and 2,436 dependency relations, the consumed time is 39 seconds. We further calculated the consumed time of each module and identified that the most time-consuming step is due to the Antlr v4 module, which is responsible for parsing files into concrete entities. In future work, we plan to investigate this bottleneck to enhance overall performance.
4. Conclusions
This tool paper proposes a prototype, Depends-Kotlin, which extracts entities and dependency relations in Kotlin code and between Kotlin and Java code. A comparison with a confirmed benchmark demonstrates Depends-Kotlin’s accuracy in capturing dependency relations and invocations in Kotlin(-Java) code. With cross-language development gaining popularity (Li et al., 2021; Mathew and Stolee, 2021; El Arnaoty and Servant, 2024), the analysis of links among different languages is essential for such software systems. To our knowledge, our prototype is the first to detect Kotlin-Java dependency relations based on static code syntax analysis. Our prototype can assist developers and researchers in handling migration process from Java to Kotlin and vice versa. Furthermore, the plug-in architecture of our prototype allows the tool to be extendable to other cross-languages, such as Kotlin-JavaScript, Kotlin-Java-JavaScript, etc. This can also aid in understanding the software systems with the interactions between different programming languages.
The results of the consumed time from three Kotlin(-Java) projects show that Depends-Kotlin can achieve good performance on small or medium-sized projects but not very well on large projects. Our future work plans to focus on improving its performance on large-scale projects.
Appendix A Walkthrough
A.1. Download and Build
Depends-Kotlin can be built and run by following 3 steps.
Step 1.:
Clone repository
Our repository is hosted on GitHub, you can clone the source code with the following command:
git clone https://github.com/XYZboom/depends-kotlin.git
Step 2.:
Compile project
Before compiling the project, make sure you have the Java 17 runtime ready. After that, run the following command at the root of your project to compile it:
on Windows
gradlew.bat build
on Linux
gradlew build
Meanwhile, the -x test option can be used to skip testing. At this point, Gradle Wapper will pull the necessary dependencies and compile them. If everything goes smoothly, you will see an output like this.
Figure 2. Screenshot of Depends-Kotlin’s compilation.
Step 3.:
Generate Runnable Jar
After successful compilation, you can use Gradle’s shadow plugin to package it as an executable jar using the following command:
on Windows
gradlew.bat shadowJar
on Linux
gradlew shadowJar
You will see the following output and find an executable jar package with the suffix *. all.jar in the projectRoot/build/libs folder.
Figure 3. Screenshot of Depends-Kotlin’s packaging.
A.2. Run from Command Line
Depends-Kotlin is based on the JVM platform and requires at least Java 17. The input of Depends-Kotlin is the source folder path of a particular project. The full command line arguments are as follows:
Usage: depends <lang> <src> <output> [OPTIONS]
Options:
--auto-include Auto include all paths under
the source path
-h, ---help Show this message and exit
-i, ---include The files of searching path
-d, ---dir The output directory
-f, ---format The output format
-g, ---granularity Granularity of dependency
-s, ---strip-leading-path Strip the leading path
-m, ---n-map-files Output DV8 dependency map file
-p, ---namepattern The name path separators
Arguments:
 lang                 The language of project files
 src                  The directory to be analyzed
 output               The output file name
Figure 4. Running screenshot of sqlex.
For example, if we need to analyze a Kotlin(-Java) project “sqlex”, we can execute the command as follows:
java -jar depends-kotlin.jar kotlin ./sqlex result -d ./out
Depends-Kotlin will analyze the source code in the ./sqlex folder and display the running process in the command console as marked in Figure 4. When the analysis is complete, the console will show the consumed time and the analysis results will be output to the specified folder. In this example, the output (analysis results) of the sqlex project can be found in the ./out folder.
Figure 5 displays the results of dependency relations at the file level in the output folder. The variables shows an ordered list of file names. The cells shows a random list of dependency relations and each cell displays the dependency relations between two source files. For example, the red rectangle shows entities in File 141 call and import entities in File 206.
Figure 5. Output screenshot of sqlex’s dependency results.

### v3

@@ 3.2. Performance Evaluation
Our experiments were conducted on a computer (AMD Ryzen 7 4800H @ 2.9GHz, 8GB RAM), and the performance of our analysis is shown in Table 4. The whole dependency extraction consists of four stages: Source File Parsing, Entity Extraction, Dependency Relation Extraction, and Result Output. We applied instrumentation in the program and calculated each stage’s running time. Due to our Entity Extraction process and Antlr’s Source File Parsing occurring simultaneously, it is difficult to split Source File Parsing and Entity Extraction. We leveraged the IntelliJ Profiler to assess the time ratio consumed by Source File Parsing, then multiplied this ratio by the total running time of the two stages. As shown in Table 4, for project sizes of 7k LOC and 51k LOC, the analysis time ranges from 32.0 seconds to 114.9 seconds. The main functions of this tool —Entity Extraction and Relation Resolution—consumes a reasonable amount of time. The most time consuming stage is Source File Parsing, which is handled by Antlr, taking 87.8%, 77.8% and 78.8% in DSU-Sideloader, Flap and RootEncoder’s total analysis time.
Table 4. Performance of the Four Stages in Three Subjects
DSU-Sideloader Flap RootEncoder
Source File Parsing 28.1s 28.8s 90.5s
Entity Extraction 0.3s 1.1s 3.9s
Relation Resolution 3.3s 6.5s 19.8s
Result Output 0.3s 0.6s 0.7s
Total 32.0s 37.0s 114.9s
4. Conclusions and Future Work
This paper proposes the Depends-Kotlin tool, which can extract dependency relations in Kotlin-Java projects. The evaluation results from three subjects show that Depends-Kotlin can capture Kotlin-Kotlin and Kotlin-Java dependencies accurately and efficiently. With cross-language development gaining popularity (Li et al., 2021; Mathew and Stolee, 2021; El Arnaoty and Servant, 2024), code changes in one language can easily propagate to and impact other languages. Our tool has the potential to assist developers in handling cross-language scenarios, such as Kotlin-Java code smell detection, code migration, and architecture analysis of such systems.
The current version of Depends-Kotlin has several limitations, such as the inability to resolve function calls on variables of generic types or determine the type of variables in lambda expressions (which results in a 85.0% recall), as well as a slow performance for source code parsing by Anltr. We plan to address these limitations in future work, with directions including modifying the Kotlin grammar rules and adopting other parsers with better performance.